OpenStudy (1dfanatic):

Complete the following for the slope intercept form equation: y=5x+6

myininaya (myininaya):

I don't understand what you want to do with y=5x+6

myininaya (myininaya):

y=mx+b is slope-intercept form where slope=m and y-intercept=b

OpenStudy (1dfanatic):

but would 6 be y then? i'm so confused..

myininaya (myininaya):

Are you looking for y-intercept?

OpenStudy (diamondboy):

that equation is already in slope intercept form

myininaya (myininaya):

I'm still now sure what you are looking for.. but if you can identify the m and b then you can easily tell me what the slope and/or y-intercept is y=mx+b y=5x+6 m=? b=?

OpenStudy (1dfanatic):

I think we are finding the y intercept but i have to put down a number and idk..

myininaya (myininaya):

well the y-intercept is what b is

myininaya (myininaya):

can you tell what b is

OpenStudy (1dfanatic):

6? sorry im just really confused and i know this is the easiest thing.

myininaya (myininaya):